Using NPV, should you invest in a project where the initial cash outflow is $28,700 and the cash inflow in the first year is $2,200 and "grows" at a rate of 2.4 percent thereafter? Assume cost of capital is 10.4 percent. (Round answer to 0 decimal places, e.g. 5,125.)

NPV of the project $
Should you invest in the project?

YesNo
Answer #1

Present value of cash inflows =Expected inflow/(r-g) =2200/(10.4%-2.4%) =27500

NPV =27500-28700 = -1200

No, the money should not be invested in the project.